# MST: Adaptive Multi-Scale Tokens Guided Interactive Segmentation

### 09 Jan 2024 v1

@@ Abstract
. In the field of Industrial Informatics, interactive segmentation has gained significant attention for its application in human-computer interaction and data annotation. Existing algorithms, however, face challenges in balancing the segmentation accuracy between large and small targets, often leading to an increased number of user interactions. To tackle this, a novel multi-scale token adaptation algorithm, leveraging token similarity, has been devised to enhance segmentation across varying target sizes. This algorithm utilizes a differentiable top-k tokens selection mechanism, allowing for fewer tokens to be used while maintaining efficient multi-scale token interaction. Furthermore, a contrastive loss is introduced to better discriminate between target and background tokens, improving the correctness and robustness of the tokens similar to the target. Extensive benchmarking shows that the algorithm achieves state-of-the-art (SOTA) performance compared to current methods. An interactive demo and all reproducible codes will be released at https://github.com/hahamyt/mst.
@@ I Introduction
IInteractive segmentation algorithms have become an important research direction in data annotation tasks and human-computer interaction tasks [1, 2, 3].
However, existing algorithms face performance bottlenecks due to challenges posed by target multi-scale variations. For example, in remote sensing segmentation task [4], targets like green land and water systems are typically much larger than buildings and roads, making it difficult to handle such complex target scale variations in single-scale features.
@@
To address poor accuracy caused by complex target scale variations, Liu et al. [2] used the SimpleFPN proposed by He et al. [5] to capture the scale variations of the vision transformer [6] features, which improved the segmentation accuracy. This SimpleFPN structure was adopted in many subsequent works [7, 8].
However, this strategy simply interpolated the single-scale ViT [6] features into different scales. This resulted in difficulty in dealing with different scale changes of segmentation targets, particularly the segmentation task of fine targets since it was difficult to inherently alleviate the limitations of single-scale features.
To improve the ViT feature performance in multi-scale scenarios, Fan et al. [9] proposed a multi-scale vision transformer model using pooling attention to simultaneously learn multi-scale feature representation across space-time resolution and channel dimension. This algorithm used different blocks to characterize multi-scale features, which is also adopted by Segformer [10]. To further enhance the multi-scale performance in interactive segmentation, Chen et al. [1] introduced Segformer’s four-stage blocks as image encoders for interactive segmentation, achieving better segmentation accuracy by utilizing equivalent multi-scale features. Nevertheless, such multi-scale feature representations still rely on a single-scale interpolation strategy, making it difficult to capture the multi-scale information of the target at the input stage.
@@
To address this issue, we proposed an interactive segmentation algorithm based on a multi-scale token fusion strategy, which adaptively selects important scale tokens as a query set to guide the base ViT tokens to pay more attention to important local regions. During token selection, important tokens are regarded as most similar to the positive clicked base tokens. Notably, the proposed algorithm can simultaneously select multi-scale important tokens, and choose the most important scale query set to inject as final scale information into the base ViT tokens.
Moreover, to improve the robustness of these important tokens, the contrastive loss [14] is used to enhance the quality of important tokens, thus achieving better segmentation. Unlike existing algorithms, the proposed one directly obtains multi-scale tokens at the input stage, which strengthens the model’s ability to handle multi-scale performance inputs.
@@
  • A learning algorithm for importance tokens based on contrastive loss is proposed, enhancing discrimination between positive clicked and background tokens.
  • As shown in Fig. 2, the proposed algorithm achieves state-of-the-art (SOTA) performance.
@@ II Related Work
The proposed algorithm primarily focuses on utilizing input multi-scale tokens in Vision Transformer [6]. As such, related work will be mainly discussed regarding the multi-scale vision transformer and important token selection.
@@ II-A Multi-scale Vision Transformer
Different from the approach of using multi-scale tokens at the input stage, Xie et al. [10] proposed a scaled self-attention, which down-sampled the key and value into larger size tokens using a scale ratio. This algorithm’s self-attention calculation is equivalent to the cross-attention between the base tokens and larger tokens. However, this strategy can only obtain larger tokens, and cannot produce tokens smaller than base tokens, as in [15].
@@ II-B Important Token Selection
To efficiently select important tokens, improve accuracy and computational efficiency, Wang et al. [18] proposed a score-based algorithm for selecting important tokens in video processing. The algorithm adopted a differentiable top-k selection algorithm proposed by Cordonnier et al. [19]. Although this top-k algorithm achieves equivalent performance with fewer tokens, This algorithm successfully obtained equivalent performance by using much fewer tokens, it requires a large amount of GPU memory when processing long sequences of tokens. Additionally, it only selects base tokens, resulting in poor multi-scale performance.
@@
Figure 3: The overall framework of the proposed algorithm
@@ III-B Multi-scale Tokens Selection and Interaction
Figure 4: The proposed multi-scale tokens interactive module
@@
where $\text{cos}(\cdot,\cdot)\in[0,1]$ gives the cosine similarity. More similar tokens have a higher cosine similarity. $j$ denotes the index of tokens of two different scales, including tiny (t) and large (l). $s_{t}\in\mathbb{R}^{L_{t}}$ and $s_{l}\in\mathbb{R}^{L_{l}}$ are similarity scores for the two scales.
After computing similarity scores between the multi-scale and base tokens, a differentiable top-k selection is proposed to individually select the top $k_{j}$ tokens with the highest scores. The top-k function from PyTorch is used to obtain the scores and corresponding indices of the top K tokens, as in equation (4).
@@
where $s_{j}^{k_{j}}$ and $\text{idx}_{j}^{k_{j}}\in\mathbb{R}^{k_{j}}$ denote the top-k scores of different scale tokens. Since these tokens vary in length, $k_{j}$ is generally set as 1/12 the length of the $L_{j}$.
Directly using $\text{idx}\_{j}^{k_{j}}$ for selection will impede gradient backpropagation. Therefore, this work uses $s_{j}^{k_{j}}$ to facilitate differentiation.
@@
To ensure the robustness of the network and its adaptability to multi-scale targets, the mean of the top-k scores $s_{t}^{k_{t}}$ and $s_{l}^{k_{l}}$ is calculated for two different scales. From these, the important tokens with the highest scores are selected as the basis for fine-tuning the interaction information $f_{b}$. As with formula (5), the one-hot matrix technique is incorporated for the differentiable selection of multi-scale tokens, which ensures proposer gradient propagation.
To facilitate better interaction, this paper computes the $K$ and $V$ matrices for $f_{j}^{k_{j}}$, as well as the $Q$ matrix for $f_{b}$. It utilizes Cross Attention[15] to fuse multi-scale information, as depicted in Fig. 4. Additionally, the down-sampled $f_{b}$ is combined with multi-scale features using Scaled Cross-Attention fusion, enabling rapid updates of the multi-scale features $f_{j},j\in{t,l}$.
@@ III-C Robust Token Selection based on Contrastive Loss
Ideally, the selected important tokens should belong to the target. However, in practice, it is found that many background tokens are wrongly selected, as shown in the first row of Fig. 8. These erroneous tokens can potentially contaminate the base tokens, thereby reducing the segmentation accuracy.
@@
To suppress the possibility of wrongly selecting important tokens in the background, the positive and negative tokens from $f_{t}^{k_{t}}$ and $f_{l}^{k_{l}}$ are separated in training, as in formula (6).
@@
where $j\in[t,l]$, $y_{j}$ symbolizes the ground truth corresponding to different patch sizes of tokens. The number of elements in $y_{j}$ equals the number of tokens in $f_{j}$.
After obtaining the positive and negative exemplar sets of important tokens, the present objective is to maximize the similarity between $f_{j}^{pos}$ and $f_{\text{kernel}}$, while minimizing the similarity between $f_{j}^{neg}$ and $f_{\text{kernel}}$. This ensures that the important tokens are predominantly derived from the actual regions of the target intended for segmentation.
After assigning the positive and negative labels to important tokens, the goal is to maximize the similarity between $f_{j}^{pos}$ and $f_{\text{kernel}}$, while minimizing the similarity between $f_{j}^{neg}$ and $f_{\text{kernel}}$. This optimization framework ensures that the important tokens are all derived from ground truth target regions.
@@
Within the proposed framework, the objective is for distance $d_{p}=|q-k_{j}^{+}|$ between $q$ and $k_{j}^{+}$ to be lesser than the distance $d_{n}=|q-k_{j}^{-}|$ between $q$ and $k_{j}^{-}$. It can be achieved by using the loss function shown in equation (7).
Another aim of the proposed algorithm is maximizing the margin between $d_{neg}$ to $d_{pos}$. This aids in imparting greater discriminative ability to the important tokens. Such discriminative ability can be obtained through the triplet patch loss function delineated in formula (7).
@@ IV-C Ablation Study
Figure 6: Distribution of object sizes in different interactive evaluation datasets. GrabCut, DAVIS, Berkeley, SBD, PascalVOC, and COCO_MVal are the image data of conventional scenes, and LoveDA and RICE are remote-sensing image scenes
@@
Figure 7: Distribution of NoC 90 of different algorithms on targets of different scales. The abscissa indicates the ratio of the target area to the total image area, and the ordinate indicates the number of clicks required to reach NoC 90.
@@
For example, when dealing with large-seized targets, existing algorithms such as SimpleClick, SAM-based algorithms, and FocalClick on the COCO MVal dataset require over 10 clicks on a substantial number of samples to achieve 90% accuracy when the target size is large. Even for small-sized targets, these algorithms also demand more than 15 clicks on a significant number of samples to reach the same level of accuracy.
In comparison, our proposed algorithm demonstrates superior performance in significantly reducing the NoC 90 for both of large and small objects, even under extreme scale changes on DAVIS, GrabCut, and Berkeley datasets.
@@
Figure 8: Results of multi-scale tokens selection. These tokens will be used to enhance the performance of base tokens in dealing with multi-scale variations
@@ IV-E Qualitative Result
As shown in Fig. 2, the proposed algorithm has a great advantage in one-click performance. To further verify this conclusion, we conducted a quantitative analysis of the one-click performance of the algorithm, and the experimental results are shown in Fig. 9.
Figure 9: The proposed MST module increases the influence range of clicks, enabling it to achieve better performance with just a single click
The algorithms involved in the comparison include SimpleClick, SAM, and HQ-SAM. It can be seen that in many cases, the proposed algorithm can achieve the effect of multiple clicks of other algorithms in one click. These quantitative results visually verify the conclusion that the proposed algorithm performs better in the case of one click. One possible reason is that the proposed MST module can effectively increase the influence range of a single click, thereby obtaining a larger receptive field than other algorithms, thereby improving performance.
In addition, in order to evaluate the performance of the proposed algorithm in multiple clicks, we analyzed the segment results of the proposed algorithm in multiple clicks to reach the target IOU, and also showed the failure cases of the proposed algorithm, the experimental results are shown in Fig. 10.
Figure 10: Quantitative analysis results of the proposed algorithm. The green mask denotes the ground truth, the purple mask denotes the prediction
As shown in Fig. 10, the segmentation results of the proposed algorithm are finer than ground truth in some cases. For example, in the first image, the ground-truth label for the target is very rough, but the proposed algorithm can segment the target well. In the third and fourth images, the proposed algorithm can also segment the clothes on the sofa and the chair by the dining table very well. In benchmarks, the ground truth of these objects is mixed with other objects, which makes it hard for the algorithm to obtain the 90% IOU after 1 click. Similar to these situations including the puppy in the 5th image and the food in the 6th image.
@@ IV-F Generalization Evaluation on Remote Sensing Images
In this paper, all the algorithms used in the comparison were trained on COCO-LVIS, a dataset that encompasses 100,000 training samples, which largely guarantees the generalization performance of the algorithms. However, this dataset only contains some common image types and scenes but lacks specific data like remote sensing images. This may lead to poor performance on remote sensing images of the algorithms that are trained on COCO-LVIS.
Nonetheless, the aforementioned situations can just right reflect the generalization performance of the model.
Therefore, the performance of the proposed algorithm and the SOTA algorithms is evaluated on the remote sensing dataset LoveDA [4] in this section. It is worth emphasizing that all algorithms were trained on COCO-LVIS and not on LoveDA. The experimental results are shown in Table IV.
@@
The data highlights the outstanding generalization capacity of the proposed algorithm in the field of remote sensing images. While SimpleClick-ViT-L also demonstrates good generalization performance, ranking second in the experiment. The proposed algorithm outperforms it by 4.7% and 1.8% in NoC 85 and NoC90 respectively. Additionally, it surpasses SimpleClick by an impressive 0.76% in NoF 90, indicating that the proposed algorithm has the best generalization performance. It is worth noting that the number of parameters of the SimpleClick-ViT-L is 1.93 times that of the proposed algorithm.
@@ V Conclusion
In this paper, we address the limitations of interactive segmentation in handling target scale variations by introducing a multi-scale feature interaction algorithm utilizing the Vision Transformer. The algorithm capitalizes on the user’s click data to derive the similarity relationships between tokens at varying scales, which allows for more effective feature screening and fusion, thereby enhancing the algorithm’s capability to handle multi-scale target transformations. Additionally, we propose a token selection strategy based on contrastive loss to increase the robustness of similarity computations. This strategy significantly improves the quality of token selection and fortifies the overall robustness of segmentation algorithm. To validate our proposed algorithm, we have subjected it to various tasks. The outcomes consistently demonstrate its superior performance in managing multi-scale target changes. Furthermore, we introduce a novel method for multi-scale performance evaluation and generalizability evaluation, providing evidence of the algorithm’s robustness and effectiveness.

### 03 Feb 2024 v2

@@ Abstract
Interactive segmentation has gained significant attention for its application in human-computer interaction and data annotation. To address the target scale variation issue in interactive segmentation, a novel multi-scale token adaptation algorithm is proposed. By performing top-k operations across multi-scale tokens, the computational complexity is greatly simplified while ensuring performance. To enhance the robustness of multi-scale token selection, we also propose a token learning algorithm based on contrastive loss. This algorithm can effectively improve the performance of multi-scale token adaptation. Extensive benchmarking shows that the algorithm achieves state-of-the-art (SOTA) performance, compared to current methods. An interactive demo and all reproducible codes will be released at https://github.com/hahamyt/mst.
@@ I Introduction
The interactive image segmentation algorithm, which can generate semantic masks for images with just a few clicks, has garnered widespread attention in recent years [1, 2, 3].
However, existing algorithms face performance bottlenecks due to the target scale variation issue. For example, in remote sensing segmentation task [4], targets like green land and water systems are typically much larger than buildings and roads, making it difficult to handle using single-scale features.
@@
To address this issue, Liu et al. [2] used the SimpleFPN proposed by He et al. [5] to capture the scale variations of the ViT [6] features, which improved the segmentation accuracy. This structure was also adopted in many subsequent works [7, 8]. However, this strategy simply interpolated the single-scale ViT [6] features into different scales. But it was difficult to inherently alleviate the limitations of single-scale features.
To obtain better multi-scale features, Fan et al. [9] proposed a multi-scale vision transformer model using pooling attention across space-time resolution and channel dimension. This algorithm used different blocks to characterize multi-scale features, which is also adopted by Segformer [10]. Chen et al. [1] introduced Segformer’s four-stage blocks as image encoders for interactive segmentation, achieving better accuracy by utilizing equivalent multi-scale features. However, this multi-scale feature is still a sampling of the features from the previous stage, making it difficult to directly obtain multi-scale features from the input.
@@
To address this issue, we proposed a multi-scale token fusion strategy, which adaptively selects important multi-scale tokens as a query set to guide the base ViT tokens to pay more attention to important local regions. To identify the important tokens, the system selects those multi-scale input tokens that are most similar to the base tokens which have been positively clicked.
Moreover, to improve the robustness of important token selection, the contrastive loss [14] is introduced to enhance the discriminability of tokens. Unlike existing algorithms, the proposed algorithm directly obtains multi-scale tokens at the input stage, which improves the performance in multi-scale target segmentation.
@@
  • A token learning algorithm based on contrastive loss is proposed, which enhances the discrimination between positively clicked tokens and background tokens.
@@ II Related Work
The proposed algorithm primarily focuses on utilizing input multi-scale tokens in ViT [6]. As such, related work will be mainly discussed regarding the multi-scale vision transformer and important token selection.
@@ II-A Multi-scale Vision Transformer
Different from using input stage multi-scale tokens algorithms, Xie et al. [10] proposed a scaled self-attention algorithm, which down-sampled the key and value into larger size tokens using a scale ratio. This algorithm’s self-attention calculation is equivalent to the cross-attention between the base tokens and larger tokens. However, this strategy can only obtain larger tokens, and cannot produce tokens smaller than base tokens.
@@ II-B Important Token Selection
To efficiently select important tokens, Wang et al. [18] proposed a score-based algorithm for selecting important tokens in video processing. The algorithm adopted a differentiable top-k selection algorithm proposed by Cordonnier et al. [19]. Although this algorithm successfully obtained equivalent performance by using much fewer tokens, it requires a large amount of GPU memory when processing long sequences of tokens. Additionally, it only selects base tokens, resulting in poor multi-scale performance.
@@
Figure 3: The overall framework of the proposed algorithm. The adaptive patch embedding module is adopted to extract multi-scale tokens adaptively, base tokens denotes the tokens with patch size $\times 16$, multi-scale tokens denotes the tokens with patch size $8\times 8$ and $28\times 28$. MST block is the proposed multi-scale token fusion module, more details can be found in Fig. 4. For efficient training, we utilize a random selection module, which has no impact during inference. The triplet token loss module represents our proposed contrastive loss-based token learning algorithm, more details can be found in Fig. 5. A simple FPN is adopted in this paper, and segmentation is performed using a two-layer MLP module
@@ III-B Multi-scale Tokens Selection and Interaction
Figure 4: The proposed multi-scale tokens interactive module. The ViT Block denotes the original vision transformer block. The Score Block is the proposed algorithm for selecting important tokens. The Cross Attention is a feature fusion algorithm; further details can be found in [fu2019dual]. The Scaled Cross Attention is an efficient cross-attention algorithm referenced in [16]. The Top-K operation is a PyTorch function that selects the largest $K$ values in a vector. The term clicked tokens refers to the local region that has been clicked
@@
where $\text{cos}(\cdot,\cdot)\in[-1,1]$ gives the cosine similarity. Tokens with similar directions have a greater cosine distance, while those with dissimilar directions have a smaller cosine distance. The sigmoid function can be mapped the scores to range [0,1], which can be regarded as the importance of token in selection. $j$ denotes the index of tokens of two different scales, including tiny (t) and large (l). $s_{t}\in\mathbb{R}^{L_{t}}$ and $s_{l}\in\mathbb{R}^{L_{l}}$ are similarity scores for the two scales.
After computing similarity scores between the multi-scale and base tokens, a differentiable top-k selection is proposed to individually select the the highest top $k_{j}$ score tokens, as shown in equation (4).
@@
where torch.topk()($\cdot$)( ⋅ ) is a Pytorch function, $s_{j}^{k_{j}}$ and $\text{idx}_{j}^{k_{j}}\in\mathbb{R}^{k_{j}}$ denote the top-k scores of different scale tokens. Since these tokens vary in length, $k_{j}$ is generally set as 1/12 the length of the $L_{j}$.
Directly using $\text{idx}\_{j}^{k_{j}}$ for selection will impede gradient backpropagation. Therefore, we uses $s_{j}^{k_{j}}$ to keep differentiable.
@@
To ensure the robustness of the network and its adaptability to multi-scale targets, the mean of the top-k scores $s_{t}^{k_{t}}$ and $s_{l}^{k_{l}}$ is calculated for two different scales. The scale with highest score will be selected to interact with $f_{b}$.
We use Cross Attention[15] to fuse multi-scale features, as illustrated in Fig. 4. $f_{b}$ will be down-sampled using Scaled Cross-Attention, allowing for efficient updating of multi-scale features $f_{j},j\in\{t,l\}$.
@@ III-C Robust Token Selection based on Contrastive Loss
Ideally, the selected important tokens should belong to the target. However, many background tokens will also be wrongly selected, which can potentially contaminate the base tokens, as shown in the first row of Fig. 8.
@@
During training, the positive and negative tokens in $f_{t}^{k_{t}}$ and $f_{l}^{k_{l}}$ will separated according to the formula (6).
@@
where $j\in[t,l]$, $y_{j}$ is the ground truth to different scale tokens. The number of elements in $y_{j}$ equals the number of tokens in $f_{j}$.
After obtaining the positive and negative exemplar sets of important tokens, the goal is to maximize the similarity between $f_{j}^{pos}$ and $f_{\text{kernel}}$, while minimizing the similarity between $f_{j}^{neg}$ and $f_{\text{kernel}}$. This ensures that the important tokens are predominantly derived from target object.
@@
The objective is for distance $d_{p}=|q-k_{j}^{+}|$ between $q$ and $k_{j}^{+}$ to be lesser than the distance $d_{n}=|q-k_{j}^{-}|$ between $q$ and $k_{j}^{-}$. It can be achieved by using the loss function shown in equation (7).
Another goal is maximizing the margin between $d_{neg}$ and $d_{pos}$. This aids in imparting greater discriminative ability to the important tokens. Such discriminative ability can be obtained through the triplet patch loss function in formula (7).
@@ IV-C Ablation Study
Figure 6: Distribution of object sizes in different interactive evaluation datasets. The horizontal axis represents the ratio of the target’s width to the image’s width, while the vertical axis represents the ratio of the target’s height to the image’s height. GrabCut, DAVIS, Berkeley, SBD, PascalVOC, and COCO_MVal are image datasets of conventional scenes. LoveDA and RICE are datasets of remote-sensing images, both of which contain bird’s-eye view images from satellites
@@
Figure 7: Distribution of NoC 90 of different algorithms on targets of different scales. The horizontal axis represents the ratio of the target area to the total image area, and the vertical axis represents the number of clicks (NoC) required to achieve target IOU 90%. The different colors of the colorbar indicate the number of samples at the current location, with more samples corresponding to a redder color. It can be observed that a larger NoC generally occur when the target size is smaller, and there are also many instances where a larger NoC occurs in scenarios where the target area is very large. It can be seen that our algorithm performs excellently across various target sizes
@@
Figure 8: Results of multi-scale tokens selection. These tokens will be used to enhance the performance of base tokens in dealing with multi-scale variations. Selected tokens are represented by transparent squares, while unselected tokens are represented by opaque squares. The top two rows of image results show the effects of using contrastive loss, while the bottom two rows demonstrate the effects without the use of contrastive loss. It can be observed that after applying contrastive loss, most of the important tokens can be accurately selected. However, when contrastive loss is not used, there are some incorrectly selected tokens present in the edge regions of the images.
@@ IV-E Qualitative Result
IV-E1 One Click Performance
As shown in Fig. 2, the proposed algorithm has a great advantage in one-click performance. The one-click experimental results are shown in Fig. 9.
Figure 9: The proposed MST module increases the influence range of clicks, enabling it to achieve better performance with just a single click. The algorithms compared include the current state-of-the-art SimpleClick, SAM, and HQSAM. It is evident that in many instances, the performance of a single click from the proposed algorithm surpasses the performance of multiple clicks from the other algorithms
In many cases, the proposed algorithm can achieve the target IOU in one click, which is significantly fewer clicks than required by other algorithms. One possible reason is that the proposed MST module can effectively increase the influence range of user clicks, obtaining a larger receptive field than others.
IV-E2 Multi-Clicks Performance
To evaluate multi-clicks performance, we analyzed the NoC results and failure cases, as shown in Fig. 10.
Figure 10: Quantitative analysis results of the proposed algorithm. The green mask denotes the ground truth, the purple mask denotes the prediction. These results are mainly used to analyze the effect of annotated data on algorithm performance. The first four rows are primarily used to analyze the differences between the algorithm’s predicted results and the labels during the process of segmentation in electric motors. It can be seen that although the algorithm can segment a particular target well, this target may not correspond to the label, which could lead to inaccurate assessment results. The last row of images analyzes the failure cases of the algorithm proposed in this paper, showing that the performance of the proposed algorithm still requires further improvement when dealing with very small and intricate targets
As shown in Fig. 10, the segmentation results of the proposed algorithm are finer than ground truth in some cases. For example, in the first image, the ground-truth label for the target is very rough, but the proposed algorithm can segment the target well. Some ground truth are mixed with other objects, which makes it hard for to obtain the 90% IOU after 1 click.
@@ IV-F Generalization Evaluation on Remote Sensing Images
In our study, all compared algorithms are trained on COCO-LVIS. However, COCO-LVIS’ focus on common image types and scenes means it lacks generalization in some special visual data like remote sensing images.
In this section, we test all methods on the remote sensing dataset LoveDA [4], as shown in Table IV.
@@
Experimental results indicate that the proposed algorithm exhibits the best generalization performance on remote sensing data.
@@ V Conclusion
In this paper, we introduced a multi-scale token interaction algorithm (MST) to address the issue of target scale variation in interactive segmentation. The MST algorithm used the user’s clicks to derive similarity relationships among tokens of different scales, thereby facilitating extension to other vision transformer-based algorithms. We also proposed a token selection strategy based on contrastive loss to increase the robustness of similarity computations, which significantly improved the quality of token selection. Extensive experiments showed the effectiveness of the proposed algorithm. Furthermore, we proposed a novel metric for evaluating multi-scale performance, which provided evidence of the algorithm’s robustness and effectiveness.